OpenStudy (anonymous):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many hours will it take for the population to grow to 100? Round to the nearest integer and do not include units in your answer. B(t) = 4 e0.8t

OpenStudy (anonymous):

\[100=4e^{.8t}\] solve for \(t\) in three steps

OpenStudy (anonymous):

1) divide by 4 2) write in equivalent logarithmic form 3) divide by \(.8\)

OpenStudy (anonymous):

\[100=4e^{.8t}\\ 25=e^{.8t}\\ \ln(25)=.8t\\ \frac{\ln(25)}{.8}=t\]

OpenStudy (anonymous):

Thank yooou! I got 4.02

OpenStudy (anonymous):

me too yw